What is the number of possible samples from the given: N = 7, n=3.


1
Expert's answer
2021-05-31T16:04:50-0400

Number of possible samples "= \frac{N!}{n!(N-n)!}"

"=\frac{7!}{3!(7-3)!} \\

= \frac{5 \times 6 \times 7}{2 \times 3} =35"